# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@
Differential Privacy of Cross-Attention with Provable Guarantee
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many more. Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about model providers and their users. In this work, we design a novel differential privacy (DP) data structure to address the privacy security of cross-attention with a theoretical guarantee. In detail, let $n$ be the input token length of system prompt/RAG data, $d$ be the feature dimension, $0<\alpha\leq 1$ be the relative error parameter, $R$ be the maximum value of the query and key matrices, $R_{w}$ be the maximum value of the value matrix, and $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods. Then, our data structure requires $\widetilde{O}(ndr^{2})$ memory consumption with $\widetilde{O}(nr^{2})$ initialization time complexity and $\widetilde{O}(\alpha^{-1}r^{2})$ query time complexity for a single token query. In addition, our data structure can guarantee that the process of answering user query satisfies $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer. Furthermore, our result is robust to adaptive queries in which users can intentionally attack the cross-attention system. To our knowledge, this is the first work to provide DP for cross-attention and is promising to inspire more privacy algorithm design in large generative models (LGMs).
@@ 1 Introduction
The development of Artificial Intelligence (AI) has four stages: (1) prediction AI, e.g., ResNet [54] in image classification; (2) generation AI, e.g., ChatGPT [1] in language generation; (3) autonomous agent AI, Voyager [105] autonomously plays Minecraft game [30]; (4) Artificial Generalization Intelligence (AGI). Humans have made rapid progress in generative AI, and we are excitingly heading to the third stage, the era of AI agent [75]. One prevalent application of AI agents is customized large generative models (LGMs) agents [80], e.g., AgentGPT [33], SuperAGI [36], MetaGPT [53, 47], GPT Researcher [35] and many so on. In particular, recently, Apple Inc. introduced Apple Intelligence [5], signaling the integration of LGMs into physical devices. This innovation allows devices to use personal information for real-life assistance, such as entering passport numbers when booking flights or informing users of their latest meetings. With increased AI capabilities, privacy concerns become significant, as the more personal information devices handle, the greater the potential privacy risks.
One fundamental technique used in LGMs is cross-attention [97], which is an essential module in retrieval-augmented generation (RAG) [68], system prompt, guided stable diffusion, and many so on. In RAG, to be more professional, the LGMs answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power. For system prompts, based on cross-attention, some customized long prompts, e.g., user information or concrete rules, are concatenated before user input to follow human instructions better, which are commonly used in ChatGPT [34], Claude3 [4] and other commercial LGMs.
Consequently, protecting the privacy of domain-specific data in RAG or system prompts is crucial as they contain sensitive information about users and companies. These data and prompts are the core assets of many start-ups. However, these data and prompts can be easily recovered [58], jailbroken [55], and released [57] by user adversarial attack [112], e.g., there are 1700 tokens in ChatGPT system prompts [83]. These findings highlight the critical importance of robust privacy protections in LGMs, making privacy not just essential but an urgent issue that demands immediate attention.
To fundamentally preserve cross-attention privacy, we borrow the powerful tools from differential privacy (DP) [19], which provides measurable privacy and combines with statistical machine learning seamlessly [84]. Thus, in this work, we would like to ask and answer the following question,
@@
Definition 1.1 (Softmax cross-attention).
Let $n$ and $m$ be the token length of the data and input query, respectively. Let $d$ be the feature dimension. Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Softmax Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is
@@
Note that $\mathsf{Softmax}(QK^{\top})=D^{-1}A\in\mathbb{R}^{m\times n}$ in Definition 1.1, which is the standard function used in transformers, and usually, we call it as attention matrix. Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
@@
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7). Let $l=\widetilde{O}(r)$. Then, our Algorithm 1 requires $O(lndr)$ memory with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output process of cross-attention satisfies $(\epsilon,\delta)$-DP and is robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
Our main technique in Theorem 1.2 ensures that cross-attention is differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem. We then solve the problem by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding. We build a data structure for weighted Softmax queries in Section 4.3, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 4.4. Furthermore, our error decreases as the input token length grows, diminishing both the relative and additive errors to zero.
Our contributions are as follows:
@@
To our knowledge, this is the first work to utilize DP to protect prompts in LGMs with theoretically provable guarantees. While some have explored protecting user/system prompts with DP [26, 77], they are primarily empirical and lack theoretical guarantees. Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs [9, 88, 63, 113, 74, 92], which diverges from our work. The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis. Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration in this area.
1.1 Related Work
@@
Differential Privacy in Data Structure and Attention. Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning. In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [2, 84], (2) during the adaptation stage: to protect target data [9, 88, 63, 113, 74, 92], (3) during the inference stage: to protect user/system prompts [26] and RAG data [68]. To protect training data, DP-SGD [2] uses DP optimizer to ensure data privacy, severing as the traditional baseline method. Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [113, 74, 31, 9, 92, 76, 88, 116, 63]. To protect user/system prompts, EW [26] provides a survey on both DP and non-DP methods. In the use of LGMs, prompting methods almost become a standard way for inference [91]. Given the billions of prompt interactions daily, ensuring privacy is essential [77]. We refer readers to Appendix A for more related works.
Roadmap. In Section 2, we present the preliminary of differential privacy (DP) and cross-attention. In Section 3, we present the main result of our cross-attention theorem (Theorem 3.1). In Section 4, we outline the main results of our algorithms. In Section 5, we discuss DP-related topics and potential extensions. In Section 6, we conclude our paper.
@@ 2 Preliminary
In this section, we give the preliminary of differential privacy (DP) and cross-attention. In Section 2.1, we describe the notations. In Section 2.2, we give definitions related to DP.
2.1 Notations
We use $\Pr[]$ to denote the probability. We use $\operatorname*{{\mathbb{E}}}[]$ to denote the expectation. We use $\operatorname*{\mathrm{Var}}[]$ to denote the variance. For two vectors $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{d}x_{i}y_{i}$. We use $X\subset\mathbb{R}^{d}$ and $|X|=n$ to mean the same thing as $X\in\mathbb{R}^{n\times d}$. Also, we denote $x_{i}^{\top}$ as the $i$-th row of $X$. We use $x_{i,j}$ to denote the $j$-th coordinate of $x_{i}\in\mathbb{R}^{n}$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e., $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$.
2.2 Differential Privacy Definitions
In this section, we give several definitions related to differential privacy (DP). We refer the reader to DR [20] for more background and details on DP.
@@ 2.2 Differential Privacy Definitions
For $\epsilon>0,\delta\geq 0$, a randomized algorithm $\mathcal{A}$ is $(\epsilon,\delta)$-DP, if for all $\mathcal{S}\subseteq\mathrm{Range}({\mathcal{A}})$ and for all $X,X^{\prime}$ such that $\|X-X^{\prime}\|_{1}\leq 1$:
@@
We mainly use the truncated Laplace mechanism, which has the following definitions.
Definition 2.4 (Truncated Laplace distribution).
We use $\mathrm{TLap}(\Delta,\epsilon,\delta)$ to denote the Truncated Laplace distribution with pdf proportional to $\exp(-\epsilon|z|/\Delta)$ on the region $[-B,B]$, where $B=\frac{\Delta}{\epsilon}\cdot\log(1+\frac{\exp(\epsilon)-1}{2\delta})$.
Fact 2.5 (Theorem 3 in GDGK [32]).
Let $z$ denote a $\mathrm{TLap}(\Delta,\epsilon,\delta)$ random variable. Then we have $\operatorname*{{\mathbb{E}}}[z]=0$, and
$\operatorname*{\mathrm{Var}}[z]=\frac{2\Delta^{2}}{\epsilon^{2}}(1-\delta\cdot\frac{\log^{2}(1+\frac{e^{\epsilon}-1}{2\delta})+2\log(1+\frac{e^{\epsilon}-1}{2\delta})}{e^{\epsilon}-1}).$
@@
Lemma 2.6 (Laplace mechanism, [20, 32], see Lemma 2.2 in AIMN [3]).
@@
Algorithm 1 Adaptive query data structure

### v3

# Provable Differentially Private Computation of the Cross-Attention Mechanism

@@
Provable Differentially Private Computation of the Cross-Attention Mechanism
Yekun Ke and Yingyu Liang and Zhenmei Shi and Zhao Song and Jiahao Zhang The University of Hong Kong.University of Wisconsin-Madison. magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.
Cross-attention has emerged as a cornerstone module in modern artificial intelligence, underpinning critical applications such as retrieval-augmented generation (RAG), system prompting, and guided stable diffusion. However, this is a rising concern about securing the privacy of cross-attention, as the underlying key and value matrices frequently encode sensitive data or private user information. In this work, we introduce a novel data structure designed to enforce differential privacy (DP) for cross-attention mechanisms, accompanied by provable theoretical guarantees. Specifically, letting $n$ denote the input sequence length, $d$ the feature dimension, $R$ the maximum magnitude of query and key matrices, $R_{w}$ the maximum magnitude of the value matrix, and $r,s,\epsilon_{s}$ the parameters for polynomial kernel methods, our proposed structure achieves $\widetilde{O}(ndr^{2})$ space and initialization complexity, with a query time of $\widetilde{O}(dr^{2})$ per token. Moreover, we demonstrate that our mechanism satisfies $(\epsilon,\delta)$-DP, incurring an additive error of $\widetilde{O}((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}R^{2s}R_{w}r^{2})$ and a relative error of $2\epsilon_{s}/(1-\epsilon_{s})$ with respect to the ground truth. Crucially, our framework maintains robustness against adaptive queries, ensuring security even in adversarial settings. To the best of our knowledge, this constitutes the first approach providing provable differential privacy for cross-attention, establishing a foundation for future privacy-preserving algorithms in large generative models (LGMs).
@@ 1 Introduction
The proliferation of Large Generative Models (LGMs) is rapidly extending beyond general-purpose chatbots towards highly personalized AI agents [86, 63]. This evolution is marked by the deep integration of LGMs into personal devices and enterprise systems, exemplified by recent initiatives like Apple Intelligence [3]. Such agents are designed to leverage personal and proprietary information to provide customized assistance, such as accessing private documents, managing schedules, or utilizing confidential company data. This personalization is frequently enabled by techniques like Retrieval-Augmented Generation (RAG) [53] and the use of detailed system prompts.
At the technical core of these capabilities lies the cross-attention mechanism [78]. Cross-attention is the fundamental module that allows the model to weigh the importance of and selectively focus on information from these external sources. While powerful, this mechanism also introduces a significant privacy vulnerability. The context that provides the LGM with its enhanced capabilities, including sensitive user data, proprietary business logic, and confidential documents, is directly processed by the cross-attention layers.
Given this, protecting the privacy of domain-specific data in RAG or system prompts becomes essential. These data and prompts are the core assets of many start-ups. However, these data and prompts can be easily recovered [44], jailbroken [41], and released [43] by user adversarial attack [92], e.g., there are 1700 tokens in ChatGPT system prompts [66]. These findings highlight the critical importance of robust privacy protections in LGMs, making privacy not just essential but an urgent issue that demands immediate attention. To fundamentally preserve cross-attention privacy, we borrow the powerful tools from differential privacy (DP) [16], which provides measurable privacy and combines with statistical machine learning seamlessly [68]. Thus, in this work, we would like to ask and answer the following question,
@@
Definition 1.1 (Softmax cross-attention, [78]).
Let $n$ and $m$ be the token length of the data and input query, respectively. Let $d$ be the feature dimension. Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, $R_{w}\geq 1$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Softmax Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is
@@
Note that $\mathsf{Softmax}(QK^{\top}/d)=D^{-1}A\in\mathbb{R}^{m\times n}$ in Definition 1.1, which is the standard function used in transformers, and usually, we call it as attention matrix. Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
@@
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma G.6). Then, our Algorithm 1 requires $\widetilde{O}(ndr^{2})$ total memory with $\widetilde{O}(ndr^{2})$ total initialization time and $\widetilde{O}(dr^{2})$ query time for a single token query, such that with probability $1-p_{f}$, the output process of cross-attention satisfies $(\epsilon,\delta)$-DP and is robust to adaptive query with relative error $2\epsilon_{s}/(1-\epsilon_{s})$ and additive error $\widetilde{O}((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}R^{2s}R_{w}r^{2})$.
Our main technique in Theorem 1.2 ensures that cross-attention is differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem. We then solve the problem by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding. We build a data structure for weighted Softmax queries in Section 5.3, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 5.4. Furthermore, our additive error decreases as the input token length grows, diminishing to zero. Our contributions are as follows:
@@
To our knowledge, this is the first work to utilize DP to protect prompts in LGMs with theoretically provable guarantees. While some have explored protecting user/system prompts with DP [22, 60], they are primarily empirical and lack theoretical guarantees. Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs [7, 72, 49, 93, 58, 75], which diverges from our work. The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis. Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration.
2 Related Works
Differential Privacy Guarantee Analysis.
Ever since [16] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [19, 51, 95, 68, 91]. DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [37, 20, 2, 47, 38, 27, 8, 11, 21, 14, 33, 61, 62, 42, 48, 23, 24, 46, 15]. Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [17]. For example, truncated Laplace mechanism [26] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions. Despite extensive research on differential privacy, existing studies on differential privacy have not yet proposed methods specifically designed to enhance the privacy of cross-attention mechanisms.
Differential Privacy in Data Structure and Attention.
@@ 2 Related Works
Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning. In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [1, 68], (2) during the adaptation stage: to protect target data [7, 72, 49, 93, 58, 75, 39], (3) during the inference stage: to protect user/system prompts [60, 22], RAG data [53], and style images [67]. To protect training data, DP-SGD [1] uses DP optimizer to ensure data privacy, serving as the traditional baseline method. Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [93, 58, 25, 7, 75, 59, 72, 49]. However, DP-SGD confines differential privacy to the optimizer. In contrast, we propose a novel approach that integrates DP directly into the attention mechanism, supported by strong theoretical analysis and guarantees. Given the resource-intensive nature of training LGMs, our technique offers a practical alternative for models trained with standard SGD, which lack inherent privacy guarantees. In such cases, applying DP-SGD would require retraining the models, which is computationally expensive, whereas our method avoids this additional cost.
Additionally, there are several recent efforts on differential privacy (DP) in attention mechanisms. For instance, [29] approximates the attention matrix with DP guarantees but targets general attention without provable bounds for weighted cross-attention Softmax. Similarly, [18] analyzes DP training challenges, such as attention distraction from noise, but focuses on full-model optimization rather than mechanism-specific privatization. Another relevant study, [10], explores DP pretraining for linear attention, differing from our polynomial kernel-based approach for full Softmax in cross-attention. Our framework is the first to provide provable DP guarantees tailored to cross-attention, offering high-probability accuracy and resilience to adaptive attacks.
Cross-Attention in System Prompt, RAG, Stable Diffusion and More.
Cross-attention [78], first introduced in language translation, is a widely used technique in many advanced AI systems. For example, Stable Diffusion [71, 57, 79, 84, 87] and SORA [64] employ cross-attention as a core module for a text-to-image conditional generation. This technique is also utilized by other multimodal models [56], including Imagen [74] and Diffusion Transformer [69]. In the realm of text-to-image editing, [35] analyzes and controls the cross-attention module to enable editing without requiring additional training. Furthermore, [94] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas. In addition, Retrieval Augmented Generation (RAG) [53, 9, 30], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module. Cross-attention also has other applications. [65] demonstrates that the prompt-tuning [55] task can be formulated as cross-attention, while [12] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation. Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [85], such as spatial reasoning [82], mathematical reasoning [50], in-context learning ability [76], compositional ability [89], few-shot adaptation ability [73, 90], and so on. There are some other works that used cross attention in Hopfield Models [40, 81, 34, 88, 80, 31, 32].
@@ 3 Preliminary
In this section, we describe the notations we use and give a preliminary of differential privacy (DP) and cross-attention.
Notations. We use $\Pr[]$ to denote the probability. We use $\operatorname*{{\mathbb{E}}}[]$ to denote the expectation. We use $\operatorname*{\mathrm{Var}}[]$ to denote the variance. For two vectors $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{d}x_{i}y_{i}$. We use $X\subset\mathbb{R}^{d}$ and $|X|=n$ to mean the same thing as $X\in\mathbb{R}^{n\times d}$. Also, we denote $x_{i}^{\top}$ as the $i$-th row of $X$. We use $x_{i,j}$ to denote the $j$-th coordinate of $x_{i}\in\mathbb{R}^{n}$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e., $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. We denote polynomial time complexity with respect to $n$ as $\operatorname{poly}(n)$. For a function $f$, we use $\widetilde{O}(f)$ to represent $f$ multiplied by a polylogarithmic factor, i.e., $f\cdot\operatorname{poly}(\log f)$. This notation, known as soft-$O$ or tilde notation, simplifies expressions by omitting logarithmic factors, focusing on the dominant term’s growth rate.
Differential Privacy Definitions. To move on, we give several definitions related to differential privacy (DP). We refer the reader to [17] for more background and details on DP.
@@
Two datasets $X,X^{\prime}\in[0,R]^{n\times d}$ are neighboring if they differ in exactly one row, i.e., there exists $i\in[n]$ such that $X_{i,*}\neq X^{\prime}_{i,*}$ and $X_{j,*}=X^{\prime}_{j,*}$ for all $j\neq i$.
@@
The sensitivity of a function $f:\mathbb{R}^{n\times d}\to\mathbb{R}^{n\times d^{\prime}}$ is: $\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n\times d}}\|f(X)-f(X^{\prime})\|_{1},$ where $X,X^{\prime}$ are neighboring datasets and $\|\cdot\|_{1}$ is the entry-wise $\ell_{1}$-norm.
@@
For $\epsilon>0,\delta\geq 0$, a randomized algorithm $\mathcal{A}$ is $(\epsilon,\delta)$-DP, if for all $\mathcal{S}\subseteq\mathrm{Range}({\mathcal{A}})$ and for all neighboring datasets $X,X^{\prime}$:
@@
To obtain the DP guarantee, we employ the truncated Laplace mechanism, which is a standard practice in differential privacy.
Definition 3.4 (Truncated Laplace distribution, [26]).
We use $\mathrm{TLap}(\Delta,\epsilon,\delta)$ to denote the Truncated Laplace distribution with pdf proportional to $\exp(-\epsilon|z|/\Delta)$ on the region [B,B][-B,B], where $B=\frac{\Delta}{\epsilon}\cdot\log(1+\frac{\exp(\epsilon)-1}{2\delta})$.
Fact 3.5 (Theorem 3 in [26]).
Let $z$ denote a $\mathrm{TLap}(\Delta,\epsilon,\delta)$ random variable. Then we have $\operatorname*{{\mathbb{E}}}[z]=0$ and
$~\operatorname*{\mathrm{Var}}[z]$
$=$ $~=\frac{2\Delta^{2}}{\epsilon^{2}}(1-\delta\cdot\frac{\log^{2}(1+\frac{e^{\epsilon}-1}{2\delta})+2\log(1+\frac{e^{\epsilon}-1}{2\delta})}{e^{\epsilon}-1}).$
@@
Lemma 3.6 (Laplace mechanism, [17, 26], see Lemma 2.2 in [2]).
@@
4 Main Results: Cross-Attention
In this section, we show our main result for cross-attention. Theorem 4.1 states that we can ensure the entire cross-attention module satisfies DP and is robust to adaptive queries. Our high-level idea is based on the similarity between weighted distance problem and cross-attention. For a typical weighted distance problem, we define the following: Let $w\in\mathbb{R}^{n}$ be the weights, $X\in\mathbb{R}^{n\times d}$ be the data matrix, where $x_{i}^{\top}$ is the $i$-th row of $X$ for $i\in[n]$, and let $y\in\mathbb{R}^{d}$ be the query. Suppose we need to answer $\ell_{1}$ distance query. We have
Algorithm 1 DP cross-attention algorithm
1:datastrucutre DPCrossAttention $\triangleright$ Theorem 4.1